OpenStudy (anonymous):

A random number generator draws at random with replacement from the digits 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. Find the chance that the digit 5 appears on more than 11% of the draws, if 1. 100 draws are made 2. 1000 draws are made

OpenStudy (anonymous):

Let X be the number of times the digit 5 appears out of n draws. X ~ Binomial(n,p). As n gets large, X ~ N(np,np(1-p)). Find P(X>11) for 1(use n=100) and P(X>110) for 2 (use n=1000).

OpenStudy (anonymous):

Thank you for your response drawar, I am however still confused on how to calculate it right now. Would you mind explaining it a bit more ? Thanks in advance

OpenStudy (anonymous):

Which part escapes you?

OpenStudy (anonymous):

@drawar Would you explaining it a bit more(part 1)???

OpenStudy (anonymous):

1. X~Bin(100,1/10) => X~N(10,9) => P(X>11)=P(Z>0.333)=.... (look up the answer in any normal distribution table). 2. Just replace n=100 by n=1000, so X~Bin(1000,1/10) =>....

OpenStudy (anonymous):

@drawar 0.3707 is correct????(part1)

OpenStudy (anonymous):

Don't know cause I don't have a table with me, just follow the hint and you SHOULD get the correct answer.

OpenStudy (anonymous):

What's the (given) correct answer?

OpenStudy (perl):

hello , i got a different answer

OpenStudy (anonymous):

@perl your ans is correct??

OpenStudy (perl):

hello

OpenStudy (perl):

I got 1- binomcdf( 100, .10, 11 ) = .296966899

OpenStudy (perl):

if you use normal approximation you should get after continuity correct normalcdf( 11.5, 10^99, 10, 3 )

OpenStudy (perl):

or using z score (11.5 - 10) / 3 =.5 =.30853

OpenStudy (perl):

or using z score P( Z > (11.5 - 10) / 3 =.5) =.30853

OpenStudy (anonymous):

@perl n wht about 2. 1000 draws are made

OpenStudy (gorv):

@abhi_abhi mate if u got 1st one than plzzz tell me

OpenStudy (anonymous):

1)0.297 2)0.135